OpenStudy (anonymous):

What is the simplified form of 1/x-2/x^2+x i got (x+2)/x(x-1) for my answer

OpenStudy (anonymous):

is this your question? \[\frac{{\frac{1}{x-2}}}{x^2+x}\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Your wording was a bit confusing... Does your equation look like @dpasingh or maybe like this? \[\frac{ 1 }{ x }- \frac{ 2 }{ x^2} + x\]

OpenStudy (anonymous):